### v4

@@ Appendix A Auxiliary Results
Besides, via (68),
$\mathbb{E}[G_{\ell}^{2}]\leq 2h_{\ell}^{-1}\big(\mathbb{E}[(X_{h_{\ell}}-X_{0})^{2}]+\mathbb{E}[(X_{h_{\ell-1}}-X_{0})^{2}]\big)\leq 2(1+M^{-1})\mathbb{E}[\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)]<\infty,$
so that $\sup_{\ell\geq 1}{\mathbb{E}[G_{\ell}^{2}]}<\infty$. Recalling that $g\mapstochar\rightarrow f_{g}(\xi)$ is bounded and that $\{g\mapstochar\rightarrow f_{X_{h_{\ell-1}}|G_{\ell}=g}(\xi),\\ \ell\geq 1\}$ are bounded uniformly in $\ell\geq 1$ by assumption, by the Cauchy-Schwarz and Markov inequalities,
@@
for some positive constant $C^{\prime}<\infty$. Furthermore, since $\sup_{\ell\geq 1}\mathbb{E}[G_{\ell}^{2}]<\infty$, $g\mapstochar\rightarrow f_{g}(\xi)$ is continuous and bounded and $G_{\ell}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits_{\ell\uparrow\infty}^{\mathcal{L}}}}G$, it holds that
@@
Finally, coming back to the decomposition (71), letting $\ell\uparrow\infty$ then $C\uparrow\infty$ yields
@@
Similar lines of reasonings using the tightness of $\{G_{\ell},\ell\geq 1\}$, the uniform boundedness and local uniform convergence of $\{(x,g)\mapstochar\rightarrow f_{X_{h_{\ell-1}}|G_{\ell}=g}(x),\ell\geq 1\}$ and the continuity of $(x,g)\mapstochar\rightarrow f_{g}(x)$ yield
@@
Since $X_{h}\mathrel{{\mathop{\rightarrow}\limits^{\mathbb{P}}}}X_{0}$ as $\mathcal{H}\ni h\downarrow 0$ by (58) and $\mathbb{P}(X_{0}=\xi)=0$,
$a(X_{0},X_{h}-X_{0})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathbb{P}}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (73)
Using (72),
$(X_{h_{\ell}}-\xi)^{+}$- $(X_{h_{\ell-1}}-\xi)^{+}=\mathds{1}_{X_{0}>\xi}(X_{h_{\ell}}-X_{h_{\ell-1}})$ (74)
@@
The standard CLT guarantees that both $\big\{h^{-\frac{1}{2}}_{\ell}(X_{h_{\ell}}-X_{0}),\ell\geq 1\big\}$ and $\big\{h^{-\frac{1}{2}}_{\ell}(X_{h_{\ell-1}}-X_{0}),\ell\geq 1\big\}$ are tight. Hence, by (73),
$h_{\ell}^{-\frac{1}{2}}\big(a(X_{0},X_{h_{\ell}}-X_{0})(X_{h_{\ell}}-X_{0})-a(X_{0},X_{h_{\ell-1}}-X_{0})(X_{h_{\ell-1}}-X_{0})\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{{\mathbb{P}}}}}0\quad\mbox{ as }\quad\ell\uparrow\infty.$ (75)
Recalling that $X_{0}$ is $Y$-measurable by definition, using (67), we get that conditionally on $Y$,
$h_{\ell}^{-\frac{1}{2}}\mathds{1}_{X_{0}>\xi}(X_{h_{\ell}}-X_{h_{\ell-1}})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$ (76)
Eventually, combining (74), (75) and (76) and invoking Slutsky’s theorem,
$h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi)^{+}-(X_{h_{\ell-1}}-\xi)^{+}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$
@@
According to (17), (15) and (40),
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}$ $=\xi^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\big(\xi^{h_{\ell}}_{N_{\ell}}-\xi^{h_{\ell}}_{\star}-\big(\xi^{h_{\ell-1}}_{N_{\ell}}-\xi^{h_{\ell-1}}_{\star}\big)\big)$ (77)
@@
recalling the definitions (36)–(39) and (41).
Similarly, from (18), (16) and (42),
@@
$h_{L}^{-1}\big(\xi^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})\geq\frac{8}{3}\gamma_{1}\lambda>1\quad\mbox{ if }\quad\beta=1.$ (78)
Using that $\sup_{h\in\mathcal{H}}{\mathbb{E}[|\xi^{h}_{0}-\xi^{h}_{\star}|]}<\infty$, that $\limsup_{n\uparrow\infty}\gamma_{n}^{-1}|\Pi_{1:n}|=0$ by (45), (78) and Lemma A.1(ii), and (21),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\big(\xi^{h_{\ell}}_{0}-\xi^{h_{\ell}}_{\star}-(\xi^{h_{\ell-1}}_{0}-\xi^{h_{\ell-1}}_{\star})\big)\Pi_{1:N_{\ell}}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Besides, via (36) and Lemma A.3(ii),
$\mathbb{E}\big[\big|g^{h_{\ell}}_{k}\big|\big]\leq|V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h_{\ell}}^{\prime\prime}(\xi^{h_{\ell}}_{\star})|\,\mathbb{E}[(\xi^{h_{\ell}}_{k-1}-\xi^{h_{\ell}}_{\star})^{2}]^{\frac{1}{2}}\leq Ch_{\ell}^{(\frac{1}{4}+\delta_{0})\wedge 1}\gamma_{k}^{\frac{1}{2}}.$ (79)
Finally, by (45), (78), Lemma A.1(i) and (21), distinguishing between the two cases $\delta_{0}<\frac{3}{4}\wedge\frac{2\beta-1}{4(1+\beta)}=\frac{2\beta-1}{4(1+\beta)}$ and $\delta_{0}\geq\frac{2\beta-1}{4(1+\beta)}$,
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(g^{h_{\ell}}_{k}-g^{h_{\ell-1}}_{k})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (48), (45), (78), Lemma A.1(i) and (21) that
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(r^{h_{\ell}}_{k}-r^{h_{\ell-1}}_{k})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Given that the innovations of the MLSA scheme are levelwise independent, the random variables $\big\{\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent with zero-mean. Moreover, at each level $\ell\geq 1$, $\{\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Thus, from (49), (45), (78), Lemma A.1(i) and (21),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k})\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},\ell\geq 1\big\}$ are independent and centered. Moreover, for any level $\ell\geq 1$, $\big\{\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},k\geq 1\big\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Hence, using (50), a comparison between series and integrals and (21),
@@
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (51), a comparison between series and integrals and (21),
@@
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\zeta^{h_{\ell}}_{k}-\zeta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (21) that
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)=h_{0}^{-\frac{3}{4(1+\beta)}}\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}}\big)^{\frac{1}{2\beta}}\frac{\sqrt{N_{0}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)}{\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}(L+1)}\big)^{\frac{1}{2\beta}}}.$ (80)
According to the decomposition (42),
@@
By (50) and (51),
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\theta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0,\quad\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\zeta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
@@
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\eta^{h_{0}}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\bigg(0,\frac{\mathbb{V}\mathrm{ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\bigg)\quad\mbox{ as }\quad N_{0}\uparrow\infty.$ (81)
Combining (80)–(81),
@@
-$\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\bigg(0,h_{0}^{-\frac{3}{2(1+\beta)}}\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}}\big)^{\frac{1}{\beta}}\frac{\mathbb{V}\mathrm{ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\bigg)\quad\mbox{ as }\quad L\uparrow\infty.$
@@
According to (21), $N_{1}\geq\dots\geq N_{L}$. Hence
@@
$=Ch_{L}^{-(2+\delta)}L^{\frac{\delta}{2}}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}^{2+\delta}|\Pi_{k+1:N_{\ell}}|^{2+\delta}\,\mathbb{E}\big[\big|e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}\big|^{2+\delta}\big].$ (82)
@@
Plugging the previous upper estimate into (82) then using (41), that $(2+\delta)\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1+\delta$ by (78), Lemma A.1(i) and (21),
@@
$=Ch_{L}^{-(2+\delta)(\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)})}L^{\frac{\delta}{2}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}^{2+\delta}}\sum_{k=1}^{N_{\ell}}\mathbb{E}\big[\big|\eta^{h_{\ell}}_{k}-\eta^{h_{\ell-1}}_{k}\big|^{2+\delta}\big].$ (84)
@@
where we used that $\mapstochar\rightarrow x^{+}, xx\in\mathbb{R}$, is $1$-Lipschitz and Lemmas A.2(i) and A.3(i). Therefore, plugging the previous upper estimate into (84) and using (21), we obtain
@@
where we used (21) and (52). Recalling the definitions (54) and (55),

### v5

@@ Appendix A Auxiliary Results
Besides, via (D.2),
$\mathbb{E}[G_{\ell}^{2}]\leq 2h_{\ell}^{-1}\big(\mathbb{E}[(X_{h_{\ell}}-X_{0})^{2}]+\mathbb{E}[(X_{h_{\ell-1}}-X_{0})^{2}]\big)\leq 2(1+M^{-1})\mathbb{E}[\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)]<\infty,$
so that $\sup_{\ell\geq 1}{\mathbb{E}[G_{\ell}^{2}]}<\infty$. Recalling that $g\mapsto f_{g}(\xi)$ is bounded and that $\{g\mapsto f_{X_{h_{\ell-1}}|G_{\ell}=g}(\xi),\\ \ell\geq 1\}$ are bounded uniformly in $\ell\geq 1$ by assumption, by the Cauchy-Schwarz and Markov inequalities,
@@
for some positive constant $C^{\prime}<\infty$. Furthermore, since $\sup_{\ell\geq 1}\mathbb{E}[G_{\ell}^{2}]<\infty$, $g\mapsto f_{g}(\xi)$ is continuous and bounded and $G_{\ell}\mathrel{{\mathop{\longrightarrow}\limits_{\ell\uparrow\infty}^{\mathcal{L}}}}G$, it holds that
@@
Finally, coming back to the decomposition (D.5), letting $\ell\uparrow\infty$ then $C\uparrow\infty$ yields
@@
Similar lines of reasonings using the tightness of $\{G_{\ell},\ell\geq 1\}$, the uniform boundedness and local uniform convergence of $\{(x,g)\mapsto f_{X_{h_{\ell-1}}|G_{\ell}=g}(x),\ell\geq 1\}$ and the continuity of $(x,g)\mapsto f_{g}(x)$ yield
@@
Since $X_{h}\mathrel{{\mathop{\to}\limits^{\mathbb{P}}}}X_{0}$ as $\mathcal{H}\ni h\downarrow 0$ by (B.24) and $\mathbb{P}(X_{0}=\xi)=0$,
$a(X_{0},X_{h}-X_{0})\mathrel{{\mathop{\longrightarrow}\limits^{\mathbb{P}}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$ (D.7)
Using (D.6),
$(X_{h_{\ell}}-\xi)^{+}$- $(X_{h_{\ell-1}}-\xi)^{+}=\mathds{1}_{X_{0}>\xi}(X_{h_{\ell}}-X_{h_{\ell-1}})$ (D.8)
@@
The standard CLT guarantees that both $\big\{h^{-\frac{1}{2}}_{\ell}(X_{h_{\ell}}-X_{0}),\ell\geq 1\big\}$ and $\big\{h^{-\frac{1}{2}}_{\ell}(X_{h_{\ell-1}}-X_{0}),\ell\geq 1\big\}$ are tight. Hence, by (D.7),
$h_{\ell}^{-\frac{1}{2}}\big(a(X_{0},X_{h_{\ell}}-X_{0})(X_{h_{\ell}}-X_{0})-a(X_{0},X_{h_{\ell-1}}-X_{0})(X_{h_{\ell-1}}-X_{0})\big)\mathrel{{\mathop{\longrightarrow}\limits^{{\mathbb{P}}}}}0\quad\mbox{ as }\quad\ell\uparrow\infty.$ (D.9)
Recalling that $X_{0}$ is $Y$-measurable by definition, using (D.1), we get that conditionally on $Y$,
$h_{\ell}^{-\frac{1}{2}}\mathds{1}_{X_{0}>\xi}(X_{h_{\ell}}-X_{h_{\ell-1}})\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$ (D.10)
Eventually, combining (D.8), (D.9) and (D.10) and invoking Slutsky’s theorem,
$h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi)^{+}-(X_{h_{\ell-1}}-\xi)^{+}\big)\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$
@@
According to (3.3), (3.1) and (B.6),
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}$ $=\xi^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\big(\xi^{h_{\ell}}_{N_{\ell}}-\xi^{h_{\ell}}_{\star}-\big(\xi^{h_{\ell-1}}_{N_{\ell}}-\xi^{h_{\ell-1}}_{\star}\big)\big)$ (E.1)
@@
recalling the definitions (B.2)–(B.5) and (B.7).
Similarly, from (3.4), (3.2) and (B.8),
@@
$h_{L}^{-1}\big(\xi^{h_{0}}_{N_{0}}-\xi^{h_{0}}_{\star}\big)\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})\geq\frac{8}{3}\gamma_{1}\lambda>1\quad\mbox{ if }\quad\beta=1.$ (E.2)
Using that $\sup_{h\in\mathcal{H}}{\mathbb{E}[|\xi^{h}_{0}-\xi^{h}_{\star}|]}<\infty$, that $\limsup_{n\uparrow\infty}\gamma_{n}^{-1}|\Pi_{1:n}|=0$ by (B.11), (E.2) and Lemma A.1(ii), and (3.7),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\big(\xi^{h_{\ell}}_{0}-\xi^{h_{\ell}}_{\star}-(\xi^{h_{\ell-1}}_{0}-\xi^{h_{\ell-1}}_{\star})\big)\Pi_{1:N_{\ell}}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Besides, via (B.2) and Lemma A.3(ii),
$\mathbb{E}\big[\big|g^{h_{\ell}}_{k}\big|\big]\leq|V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h_{\ell}}^{\prime\prime}(\xi^{h_{\ell}}_{\star})|\,\mathbb{E}[(\xi^{h_{\ell}}_{k-1}-\xi^{h_{\ell}}_{\star})^{2}]^{\frac{1}{2}}\leq Ch_{\ell}^{(\frac{1}{4}+\delta_{0})\wedge 1}\gamma_{k}^{\frac{1}{2}}.$ (E.3)
Finally, by (B.11), (E.2), Lemma A.1(i) and (3.7), distinguishing between the two cases $\delta_{0}<\frac{3}{4}\wedge\frac{2\beta-1}{4(1+\beta)}=\frac{2\beta-1}{4(1+\beta)}$ and $\delta_{0}\geq\frac{2\beta-1}{4(1+\beta)}$,
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(g^{h_{\ell}}_{k}-g^{h_{\ell-1}}_{k})\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (B.14), (B.11), (E.2), Lemma A.1(i) and (3.7) that
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(r^{h_{\ell}}_{k}-r^{h_{\ell-1}}_{k})\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Given that the innovations of the MLSA scheme are levelwise independent, the random variables $\big\{\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k}),\ell\geq 1\big\}$ are independent with zero-mean. Moreover, at each level $\ell\geq 1$, $\{\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Thus, from (B.15), (B.11), (E.2), Lemma A.1(i) and (3.7),
@@
$h_{L}^{-1}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}\Pi_{k+1:N_{\ell}}(\rho^{h_{\ell}}_{k}-\rho^{h_{\ell-1}}_{k})\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
$\big\{\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},\ell\geq 1\big\}$ are independent and centered. Moreover, for any level $\ell\geq 1$, $\big\{\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k},k\geq 1\big\}$ are $\{\mathcal{F}^{h_{\ell}}_{k},k\geq 1\}$-martingale increments. Hence, using (B.16), a comparison between series and integrals and (3.7),
@@
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\theta^{h_{\ell}}_{k}-\theta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
Using (B.17), a comparison between series and integrals and (3.7),
@@
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}}\sum_{k=1}^{N_{\ell}}\zeta^{h_{\ell}}_{k}-\zeta^{h_{\ell-1}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad L\uparrow\infty.$
@@
It follows from (3.7) that
$h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)=h_{0}^{-\frac{3}{4(1+\beta)}}\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}}\big)^{\frac{1}{2\beta}}\frac{\sqrt{N_{0}}\big(\chi^{h_{0}}_{N_{0}}-\chi^{h_{0}}_{\star}\big)}{\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}(L+1)}\big)^{\frac{1}{2\beta}}}.$ (E.4)
According to the decomposition (B.8),
@@
By (B.16) and (B.17),
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\theta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0,\quad\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\zeta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad N_{0}\uparrow\infty.$
@@
$\frac{1}{\sqrt{N_{0}}}\sum_{k=1}^{N_{0}}\eta^{h_{0}}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\bigg(0,\frac{\operatorname{\mathbb{V}ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\bigg)\quad\mbox{ as }\quad N_{0}\uparrow\infty.$ (E.5)
Combining (E.4)–(E.5),
@@
$\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\bigg(0,h_{0}^{-\frac{3}{2(1+\beta)}}\big(1-M^{-\frac{2\beta-1}{2(1+\beta)}}\big)^{\frac{1}{\beta}}\frac{\operatorname{\mathbb{V}ar}\big((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+}\big)}{(1-\alpha)^{2}}\bigg)\quad\mbox{ as }\quad L\uparrow\infty.$
@@
According to (3.7), $N_{1}\geq\dots\geq N_{L}$. Hence
@@
$=Ch_{L}^{-(2+\delta)}L^{\frac{\delta}{2}}\sum_{\ell=1}^{L}\sum_{k=1}^{N_{\ell}}\gamma_{k}^{2+\delta}|\Pi_{k+1:N_{\ell}}|^{2+\delta}\,\mathbb{E}\big[\big|e^{h_{\ell}}_{k}-e^{h_{\ell-1}}_{k}\big|^{2+\delta}\big].$ (E.6)
@@
Plugging the previous upper estimate into (E.6) then using (B.7), that $(2+\delta)\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1+\delta$ by (E.2), Lemma A.1(i) and (3.7),
@@
$=Ch_{L}^{-(2+\delta)(\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)})}L^{\frac{\delta}{2}}\sum_{\ell=1}^{L}\frac{1}{N_{\ell}^{2+\delta}}\sum_{k=1}^{N_{\ell}}\mathbb{E}\big[\big|\eta^{h_{\ell}}_{k}-\eta^{h_{\ell-1}}_{k}\big|^{2+\delta}\big].$ (E.8)
@@
where we used that $\mapsto x^{+}, xx\in\mathbb{R}$, is $1$-Lipschitz and Lemmas A.2(i) and A.3(i). Therefore, plugging the previous upper estimate into (E.8) and using (3.7), we obtain
@@
where we used (3.7) and (B.18). Recalling the definitions (B.20) and (B.21),